# Formalising the intentional stance: attributing goals and beliefs to stochastic processes

### v1

@@ 1.1 Outline of our approach and results
Finally, we touch on the issue of bounded rationality. A transducer may fail to be be optimal for a given teleo-environment $\varepsilon$ while still being optimal for $\varepsilon$ within some constrained class of transducers, meaning simply a subset of all transducers. These constrained classes of transducers are intended to model limitations in the cognitive capacity of a type of agent, such as a limited supply of memory, or of randomness. For some kinds of limitation (i.e. some constrained classes of transducers), the value-laden Bellman property still holds. (In particular, it holds for constrained classes obeying the so-called exchange property, Definition 3.3.1.) For agents that are optimal within such a class it is permissible to attribute normative-epistemic states that update by value-laden filtering, as described above. However, this is not true for all constrained classes of interest. We demonstrate this by applying our framework to a version of the ‘absent-minded driver,’ [32] a classic problem from decision theory that involves making decisions without memory. The optimal solution to the absent-minded driver problem is stochastic rather than deterministic, and as a transducer it is not possible to give it a mental state attribution that updates by value-laden filtering.
1.2 Formal methodology: the coinductive approach
Our formal treatment uses coinductive reasoning to construct the objects we call transducers. While the coinductive formalism might not be familiar to most readers, it corresponds closely to the intuition of ‘moving step-by-step’ and produces elegant proofs. For any readers wishing to become familiar with the formal details of coinductive reasoning we recommend [9] for a set-theoretic justification for the approach we use in this paper, which treats coinduction as a first-class primitive. Alternatively, [22] provides a comprehensive introduction from the perspective of category theory, and [1] a type-theoretic approach.
Intuitively, while induction lets us define and reason about objects with a finite internal structure, such as natural numbers or finite trees, coinduction lets us define and reason in a similar way about objects with an infinite internal structure. In practice this amounts to accepting objects that do not start at an inductive base case, such as trees without leaf nodes, which are necessarily infinite. Coinduction is the category-theoretic dual of mathematical induction, and therefore many of the concepts associated with induction have coinductive analogs. The main mathematical consequence of this is that one can define functions into coinductive objects using the function being defined [9, Theorem 2.4]; [1, Theorem 7] and prove theorems about relations that are coinductively defined [9, Theorem 4.29] [^2] No theorem for this is needed in [1] due to type theory treating propositions as types. or equality [9, Theorem 4.33][1, Theorem 18] using coinductive hypotheses (by analogy to inductive hypotheses).
Note that we could have derived the same results without using coinduction, but we chose to use the coinductive formalism in order to emphasise the dynamical relations between transducers, permitting an elegant set of definitions and proofs.
2 Relation To Previous Work
This idea can be seen as an attempt to formalise Dennett’s intentional stance [10, 12, 13]; in doing so, we are also able to formally model relationships between (analogues of) Dennett’s intentional stance and his physical stance. In the current work, the physical stance corresponds roughly to our treatment of stochastic Moore machines and unifilar machines (Section 3.3), while the intentional stance correpsonds to our exploration of the relationship between transducers and teleo-environment, explored in Section 4 and throughout the paper.
In the past there have been significant disagreement about what sorts of system should be regarded as ‘agents’. Some authors (e.g. [34, p.34] or [43]) categorise even simple feedback systems (such as thermostats or guided torpedoes) as agents. Others (e.g. [14] or [23]) explicitly reject such a characterisation, claiming that thermostats may behave (in some ways) as if they were agents, but are not truly agents. A discussion of this general debate can be found in [25]. Currently, echoes of this disagreement can be seen in disputes about the extent to which AI systems such as large language models (LLMs) should be understood as agents. The idea of ‘as-if’ agency is that we should not ignore simple systems such as thermostats when studying agency, even if the skeptics are correct and thermostats lack some important property. It is still valuable to have a theory that explains in what sense even a thermostat behaves as though it were an agent.
We will not define what a ‘formal theory of rationality’ is, but our notion of teleo-environment (Section 4) will serve as an example.
In the current paper we are also concerned with consistency. We show that we obtain a form of consistency (Section 5) if certain assumptions are met. The form of consistency is not what we might naively expect, since the update rule is given by ‘value-laden filtering’ (Definition 5.0.1), rather than vanilla Bayesian filtering. We also observe that this ‘value-laden filtering’ consistency does not hold in general in bounded-rational contexts. In particular, in Section 5.1, we describe a case in which it doesn’t hold, due to constraints on memory. In Section 6.5 we also discuss another consequence of optimality under memory constraints, namely that the uniquely optimal policy may be non-deterministic, as previously observed in the absent-minded driver problem ([32]) and discussed more generally in [21].
2.1 Computational mechanics and unifilar machines
On a technical level our work is related to the tools of computational mechanics [36]. As in computational mechanics we are concerned with the externally observable behaviour of stochastic input-output processes, and the way in which we do this has something in common with the idea of an $\varepsilon$-transducer [3]. The formal objects known as $\varepsilon$-machines and $\varepsilon$-transducers are the minimal unifilar representations of stochastic processes (usually assumed to be stationary), and their states are called causal states. The concept of unifilarity plays an important role in our work (see Section 3.3.2). This concept originally comes from the computational mechanics literature, since it is used in one of the ways that one can define an $\varepsilon$-machine [38], or more generally an $\varepsilon$-transducer [3]. However, in our case, instead of using a minimal unifilar representation we are concerned with the equivalent of causal states as mathematical objects in their own right (our transducers). The connection between these ideas is made more precise in [39], where transducers are shown to form the states of a single unifilar machine. (The terminal one, in the sense of category theory.) We also note that the mapping from stochastic Moore machines to unifilar machines (our definition 3.3.7) has much in common with the ‘mixed state representation’ of a stochastic processes described in [8].

### v2

@@ 3.5 Bounded rationality
Finally, we touch on the issue of bounded rationality. A transducer may fail to be be optimal for a given teleo-environment $\varepsilon$ while still being optimal for $\varepsilon$ within some constrained class of transducers, meaning simply a subset of all transducers. These constrained classes of transducers are intended to model limitations in the cognitive capacity of a type of agent, such as a limited supply of memory, or of randomness. For some kinds of limitation (i.e. some constrained classes of transducers), the value-laden Bellman property still holds.
In particular, the value-laden Bellman property holds for constrained classes that are ‘closed under trajectory splicing’, which is a somewhat technical property given in Definition 3.3.3 in the companion paper. For the interested reader, here is a brief informal description: given transducers $\pi,\pi^{\prime}$ and a sequence of inputs and outputs $(\mathbf{i},\mathbf{o})$, define a new transducer called “$\pi$ spliced with $\pi^{\prime}$ along $(\mathbf{i},\mathbf{o})$”, which behaves exactly like $\pi$, except that if it happens to produce the sequence o in response to the input sequence i starts behaving exactly as $\pi^{\prime}$ would have done in the same situation. Then a constrained class of transducers $T$ is closed under trajectory splicing if whenever $\pi,\pi\in T$, we also have that $\pi$ spliced by $\pi^{\prime}$ along $(\mathbf{i},\mathbf{o})$ is in $T$, whenever $(\mathbf{i},\mathbf{o})$ is a possible sequence of inputs and outputs.
For agents that are optimal within such a class it is permissible to attribute normative-epistemic states that update by value-laden filtering, as described above. However, this is not true for all constrained classes of interest. We demonstrate this in Section 6.5 of the companion paper by applying our framework to a version of the ‘absent-minded driver,’ [25] a classic problem from decision theory that involves making decisions without memory. The optimal solution to the absent-minded driver problem is stochastic rather than deterministic, and as a transducer it is not possible to give it a mental state attribution that updates by value-laden filtering.
4 Relation To Other Frameworks
Our approach has some parallels with existing bodies of work in decision theory, cognitive neuroscience, machine learning and complex systems. In this section we discuss the relation of our approach to representation theorems in decision theory, the free-energy framework, inverse reinforcement learning, and computational mechanics.